A ball is thrown vertically upwards with a velocity u from the baloo descending with velocity V. The ball will pass by the ball after time,


a) u - V/2g


b) u+V/2g


c) 2(u-V)/g


d) 2(u+V)/g

Grade:11

1 Answers

raj mishra
11 years ago

ans d= 2(u+V)/g;

Vball wrt balloo = u+V

Srelative= (u+V)t - 1/2gtsquare=0

therefore t= opt d